# SwiftVideo: A Unified Framework for Few-Step Video Generation through Trajectory-Distribution Alignment

### v1

@@ Abstract
Introduction
Recent advancements in diffusion and flow-based models have significantly enhanced generative modeling, especially in video synthesis. However, these models typically require many iterative sampling steps, resulting in high computational and time costs that limit their practical deployment. Although higher-order numerical ODE solvers (Lu et al. 2025, 2022) can reduce the number of sampling steps, they often suffer from severe quality degradation when operating in the few-step regime (i.e., $\textnormal{steps}\leq 10$).
To address these challenges, diffusion distillation methods have emerged to reduce sampling overhead in video generation. Trajectory-preserving approaches, such as LCM, PCM (Luo et al. 2023; Wang et al. 2024), leverage consistency loss to minimize inference steps. However, as discrete-time consistency models (CMs), they tend to produce blurry outputs under few-step settings and introduce additional discretization errors from numerical ODE solvers. Alternatively, distribution-matching methods like DMD2 (Yin et al. 2024a), train student models to approximate the teacher’s distribution in order to reduce inference steps. However, this approach inherently upper-bounds the student model’s performance by the teacher model and can lead to domain inconsistencies due to relying solely on distributional constraints. Other recent strategies (Ren et al. 2024; Sauer et al. 2024a) incorporate post-training enhancements for few-step generation, but often depend on auxiliary reward models or expensive-to-collect preference datasets.
Figure 1: Video generation results on the OpenVid-1M dataset (Nan et al. 2025) using 4 sampling steps. Red circles and boxes highlight blurry regions and artifacts in the generated videos. DA and TA denote Distribution Alignment and Trajectory Alignment, respectively. SwiftVideo w/o DA & TA employs only Continuous-time Consistency Distillation (CCD). Our method significantly outperforms PCM.
@@ Introduction
Firstly, we introduce CCD to ensure precise ODE trajectory preservation, eliminating the discretization errors that arise from ODE solver reliance in discrete-time consistency models. As shown in Figure 1, while CCD alone outperforms PCM, it still exhibits blurriness. Nevertheless, its accurate trajectory preservation enables effective subsequent training optimization. Secondly, instead of approximating the teacher’s distribution, we directly perform DA on real data, thereby breaking the performance constraints imposed by the teacher model. Thirdly, we observe that the distribution-aligned distilled model exhibits significant improvement in generation quality as the number of steps increases within the low-step regime (i.e., $\text{steps}\leq 8$). Based on this observation, we propose TA across different inference steps as a post-training strategy. Specifically, we generate the preference dataset using the distilled model with varying numbers of steps, then leverage the Direct Preference Optimization (DPO) algorithm (Rafailov et al. 2023) to implicitly align low-step inference trajectories with high-step inference trajectories, eliminating the need for reward models and costly external preference data collection. As shown in Figure 1, DA effectively reduces blurriness while TA improves detail preservation under few-step settings.
@@
Related Work
Video Diffusion Model
Driven by advances in large-scale video data and model architectures, video diffusion models have achieved remarkable success. Early approaches (Blattmann et al. 2023; Guo et al. 2024) using 3D U-Net architectures in latent space were limited to generating short, low-quality videos. With the success of Sora (OpenAI 2024), modern models have shifted toward Diffusion Transformer (DiT) architectures, with open-source implementations like HunyuanVideo (Kong et al. 2025), CongVideoX (Yang et al. 2025), and Wan (Wan et al. 2025). Although these models can produce longer, higher-quality videos, generation remains computationally expensive and time-consuming. To address this limitation, we leverage diffusion distillation methods to accelerate these models without compromising their quality.
Diffusion Model Distillation
@@
Trajectory-preserving distillation aims to train a student model capable of generating samples in fewer steps by approximating the ordinary differential equation (ODE) sampling trajectory of the teacher model. Progressive Distillation (Salimans and Ho 2022) trains the student model to predict the subsequent flow locations determined by the teacher model. Rectified Flow (Liu, Gong, and Liu 2022) trains the student model on noisy-image pairs obtained from the teacher, thereby progressively straightening the ODE trajectory. Consistency Distillation (Luo et al. 2023) trains a student model to map noisy samples along the ODE trajectory to its origin. Although trajectory-preserving distillation facilitates training, the samples generated in few-step tend to exhibit blurriness and inferior quality.
Distribution-matching distillation trains a student model to approximate the teacher’s distribution. Some methods leverage adversarial loss functions to improve the generative quality of student models (Sauer et al. 2024b, a). DMD (Yin et al. 2024b) minimizes approximate KL divergence using gradients from two score functions: a fixed pretrained target score and a dynamically trained generator score.
Recently, trajectory-preserving distillation methods (Li et al. 2024; Wang et al. 2023) and distribution-matching distillation(Zhang et al. 2024)) have been increasingly applied to video generative models. However, many of these distilled video models originate from the pre-Sora era and typically generate only low-quality, short-duration results. Moreover, existing consistency distillation methods for video are trained using discretized timesteps, which rely on numerical ODE solvers and introduce additional discretization errors. In contrast, our approach employs continuous-time consistency distillation to address these issues. Furthermore, prior distribution-matching distillation methods usually approximate the teacher’s output distribution as the target, which implicitly limits the generation quality by the teacher model. Instead, our approach directly performs distribution alignment on real data, similar to (Lin et al. 2025).
@@
Preliminary
Flow Matching
Flow Matching (Lipman et al. 2023; Liu, Gong, and Liu 2022) is widely used in recent advanced video-generation models (Wan et al. 2025). Given a data sample $\mathbf{x}_{0}\sim p_{\textnormal{data}}$ and a noise sample $\mathbf{x}_{1}\sim p_{\textnormal{noise}}$, the noisy data $\mathbf{x}_{t}$ is defined as a linear interpolation between these two distributions:
@@
where $\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\mathbf{x}_{t},t)}{\mathrm{d}t}=\nabla_{x_{t}}\bm{f}_{\theta^{-}}(\mathbf{x}_{t},t)\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}+\partial_{t}\bm{f}_{\theta^{-}}(\mathbf{x}_{t},t)$ is the tangent of $\bm{f}_{\theta^{-}}$ at $(\mathbf{x}_{t},t)$ along the trajectory of the PF-ODE $\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$. Continuous-time CMs do not rely on ODE solvers, which avoids discretization errors and offers more accurate supervision signals during training. Recently, sCM (Lu and Song 2025) presents several training tricks to simplify, stabilize, and scale the training of continuous-time consistency models, achieving superior performance compared to discrete-time consistency models.
Method
@@
Continuous-time Consistency Distillation
Algorithm 1 Continuous-time Consistency Distillation
@@
$\bm{F}_{\theta}(x,t)$ is a deep neural network. Unlike the TrigFlow formulation employed in sCM (Lu and Song 2025), we set $c_{\mathrm{skip}}=1$ and $c_{\mathrm{out}}=-t$, which aligns with the Rectified Flow (Liu, Gong, and Liu 2022) formulation commonly adopted in current video diffusion models and eliminates the need for additional formulation conversion. Moreover, this parameterization satisfies the boundary conditions and is compatible with the Euler ODE solver. Then we differentiate both sides of Eq (4) with respect to $t$:
@@
Training Stabilization and Acceleration
Building upon the approach proposed in sCM (Lu and Song 2025), we employed two training techniques, Tangent Normalization and Tangent Warmup, to mitigate the instability issues in the continuous-time consistency distillation process. Additionally, given that current video diffusion models comprise multiple DiT blocks, we designed an iterative algorithm to compute the tangent function $\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\mathbf{x}_{t},t)}{\mathrm{d}t}$ with minimal computational overhead.
@@
Alignment for Distilled Model
@@
Distribution Alignment
While continuous-time consistency distillation achieves precise ODE trajectory preservation, the generated results still exhibit blurriness due to the inherent properties of consistency models (Lin, Wang, and Yang 2024). Therefore, we introduce distribution alignment to improve the visual fidelity and realism of generated videos. Unlike previous distribution-matching methods that constrain the upper limit of the student’s generation quality by matching the teacher’s distribution, we directly employ adversarial loss to align with real data distribution.
The distribution alignment employs a discriminator trained to distinguish between generated samples $\hat{\mathbf{x}}_{0}$ and real data $\mathbf{x}_{0}$, where $\hat{\mathbf{x}}_{0}=\mathbf{x}_{t}-t\bm{F}_{\theta}(\mathbf{x}_{t},t)$. For the discriminator, we follow the design in StyleGAN-T (Sauer et al. 2023). We utilize a frozen pretrained feature network DINOv2 (Oquab et al. 2024) coupled with trainable lightweight discriminator heads, comprising both temporal heads for capturing inter-frame temporal dynamics and spatial heads for evaluating intra-frame visual quality.
@@
Subsequently, we finetune our model using Direct Preference Optimization (Liu et al. 2025) on this synthetic preference dataset. As shown in Figure 3, since the synthetic data implicitly contains flow trajectory information, this training enables the model to align with high-step inference trajectories, thereby enhancing low-step generation quality. The DPO loss $\mathcal{L}_{\textnormal{DPO}}$ on flow-based video generation model is written below:
@@
Employing Eq. (11) alone as the training objective for trajectory alignment results in undesired training dynamics. As training progresses, the model minimizes the loss by reducing velocity prediction accuracy on dispreferred data more than on preferred data, thereby causing the flow to diverge from the intended trajectory. To address this issue, we introduce a regularization term $\|\mathbf{v}^{w}-\mathbf{v}_{\theta}(\mathbf{x}_{t}^{w},t)\|^{2}$ to maintain high prediction fidelity on the preferred data. Given that our preference dataset is entirely synthetic, this regularization term takes the same form as the reflow loss (Liu, Gong, and Liu 2022), denoted as $\mathcal{L}_{RF}$. Therefore, the overall trajectory alignment loss can be formulated as $\mathcal{L}_{\textnormal{DPO}}+\lambda_{\textnormal{RF}}\cdot\mathcal{L}_{\textnormal{RF}}$, where $\lambda_{\textnormal{RF}}$ is a hyper-parameter. We note that the model performance is robust to the choice of $\lambda_{\textnormal{RF}}$, as detailed in the appendix.
Experiment
Implementation Details.
@@
In our distillation experiments, we use Wan2.1-FUN-inp-480p-1.3B (Wan et al. 2025) as the foundational model. For fair comparison, all the experiments are conducted on OpenVid-1M (Nan et al. 2025), a large-scale, high-quality video dataset that provides a standardized benchmark for few-step video generation. Specifically, we randomly select one million videos for training and 1000 for evaluation. In the continuous-time consistency distillation and distribution alignment, we fix the resolution of the training videos to $832\times 480$ with 61 frames and a batch size of 1. We conduct full-parameter training for 3,000 iterations using a learning rate of $1\mathrm{e}-6$ for both the discriminator and student model. The default ema rate is 0.95, warm-up iteration $\mathcal{N}_{\textnormal{warmup}}$ is 1000 and $\lambda_{\textnormal{adv}}$ is set to 0.01. In trajectory alignment, we first utilize distribution-aligned distilled model to generate videos using 4 steps and 8 steps sampling respectively, then construct preference dataset of size 5000. Then we apply LoRA (Hu et al. 2022) to fine-tune the model’s linear layers on this preference dataset, using a learning rate of $1\mathrm{e}-6$ for approximately 2000 iterations. To further improve 2-step generation quality, we conduct an additional round of trajectory alignment, employing the once-aligned model to generate preference data using 2-step and 4-step sampling with the same training procedure. $\lambda_{RF}$ is set to 2 and $\beta$ in $\mathcal{L}_{\mathrm{DPO}}$ is set to 2500. All training uses the AdamW optimizer (Loshchilov and Hutter 2019), and we employ the Euler solver for sampling.
Evaluation Metrics
@@
Following existing work (Zhang et al. 2024; Yin et al. 2025), we use FVD (Ge et al. 2024) and standard VBench-I2V benchmark (Huang et al. 2024) to evaluate our model. FVD measures the Fréchet distance between feature distributions of real and generated videos. VBench is a comprehensive benchmark that is widely used to evaluate the video quality. All evaluations are conducted on the OpenVid-1M test set.
Comparisons on Image-to-Video Generation
Quantitative Comparisons.
@@
We conduct comprehensive evaluations by comparing our model against state-of-the-art distillation methods, including trajectory-preserving methods LCM (Luo et al. 2023) and PCM (Wang et al. 2024), and distribution-matching methods OSV (Mao et al. 2025) and DMD2 (Yin et al. 2024a). We implement these methods on Wan 2.1 following the official codebases of LCM, PCM, and DMD2. And we try our best to implement OSV, which we denote as $\textnormal{OSV}^{\dagger}$. During evaluation, we utilize the first frame from the Openvid-1M test set as conditional input to generate 1000 videos at 832×480 resolution with 24 FPS, then compute FVD and VBench metrics on these generated videos.
@@
Figure 4: Results from our user study regarding frame quality, temporal consistency, and image alignment of different models. The teacher method uses 8 sampling steps, while other methods use 4 steps.
Qualitative Comparisons
Figure 5 presents visual comparison results of the teacher model with 8 sampling steps, LCM (Luo et al. 2023), PCM (Wang et al. 2024), OSV (Mao et al. 2025), DMD2 (Yin et al. 2024a) and our method with 4 inference steps. Figure 5 demonstrates that our method generates videos with superior visual clarity, richer details, and enhanced realism compared to other approaches, while exhibiting fewer artifacts. Additional qualitative results are provided in the appendix.
To further evaluate the performance of our method, we conduct a comprehensive user study comparing it against existing distillation approaches and the teacher model. We select first frames from 50 videos across different categories in the OpenVid-1M test set as conditional inputs. For each image, we generate 5 videos using same random seeds, with distillation methods using 4 sampling steps and the teacher model using 8 steps. Participants evaluate and select the best video based on frame quality, temporal consistency, and image alignment. As shown in Figure 4, our approach significantly outperforms baseline distillation methods and the teacher model with 8 sampling steps.
Figure 5: Qualitative comparison results. Our results significantly outperform other methods in both frame quality and temporal consistency under 4 sampling steps. All visual results are generated using images from the OpenVid-1M (Nan et al. 2025) test set.
Ablation Studies
Effect of Continuous-time Consistency Distillation (CCD)
To validate continuous-time consistency distillation effectiveness, we evaluate three training strategies: direct distribution alignment (DA), discrete-time consistency distillation (DCD) followed by DA, and continuous-time consistency distillation (CCD) followed by DA. Table 2a demonstrates that consistency distillation combined with distribution alignment achieves superior performance. DA alone produces excessive trajectory deviation from the original model. Furthermore, CCD outperforms DCD through more accurate ODE trajectory preservation.
Figure 6: Effect of synthetic preference dataset and reflow loss in trajectory alignment.
Effect of Distribution Alignment (DA)
To validate the effectiveness of distribution alignment, we conduct ablation studies using only continuous-time consistency loss for distillation. As shown in Table 2b, DA yields superior performance by directly approximating the real data distribution, which enhances visual fidelity under few-step sampling.
Effect of Trajectory Alignment (TA)
Comparing distilled models with and without trajectory alignment, Table 2c demonstrates that model refined through TA generated higher-quality videos. This improvement stems from using DPO to align the model with high-step inference trajectories, thereby enhancing low-step inference quality.
Time Samplers
Prior work (Esser et al. 2024) demonstrates that the distribution of timestep affects generation quality. We train models using CCD with DA under different timestep distributions. As shown in Table 2d, the logit-normal timestep sampler achieves superior results.

### v2

@@ Abstract
1 Introduction
footnotetext: $\dagger$ Equal contribution.
Recent advancements in diffusion and flow-based models have significantly enhanced generative modeling, especially in video synthesis. However, these models typically require many iterative sampling steps, resulting in high computational and time costs that limit their practical deployment. Although higher-order numerical ODE solvers (Lu et al., 2025, 2022) can reduce the number of sampling steps, they often suffer from severe quality degradation when operating in the few-step regime (i.e., $\textnormal{steps}\leq 10$).
To address these challenges, diffusion distillation methods have emerged to reduce sampling overhead in video generation. Trajectory-preserving approaches, such as LCM, PCM (Luo et al., 2023; Wang et al., 2024), leverage consistency loss to minimize inference steps. However, as discrete-time consistency models (CMs), they tend to produce blurry outputs under few-step settings and introduce additional discretization errors from numerical ODE solvers. Alternatively, distribution-matching methods like DMD2 (Yin et al., 2024a), train student models to approximate the teacher’s distribution in order to reduce inference steps. However, this approach inherently upper-bounds the student model’s performance by the teacher model and can lead to domain inconsistencies due to relying solely on distributional constraints. Other recent strategies (Ren et al., 2024; Sauer et al., 2024a) incorporate post-training enhancements for few-step generation, but often depend on auxiliary reward models or expensive-to-collect preference datasets.
@@ 1 Introduction
Figure 1: Video generation results on the OpenVid-1M dataset (Nan et al., 2025) using 4 sampling steps. Red circles and boxes highlight blurry regions and artifacts in the generated videos. DA and TA denote Distribution Alignment and Trajectory Alignment, respectively. SwiftVideo w/o DA & TA employs only Continuous-time Consistency Distillation (CCD). Our method significantly outperforms PCM.
Firstly, we introduce CCD to ensure precise ODE trajectory preservation, eliminating the discretization errors that arise from ODE solver reliance in discrete-time consistency models. As shown in Figure 1, while CCD alone outperforms PCM, it still exhibits blurriness. Nevertheless, its accurate trajectory preservation enables effective subsequent training optimization. Secondly, instead of approximating the teacher’s distribution, we directly perform DA on real data, thereby breaking the performance constraints imposed by the teacher model. Thirdly, we observe that the distribution-aligned distilled model exhibits significant improvement in generation quality as the number of steps increases within the low-step regime (i.e., $\text{steps}\leq 8$). Based on this observation, we propose TA across different inference steps as a post-training strategy. Specifically, we generate the preference dataset using the distilled model with varying numbers of steps, then leverage the Direct Preference Optimization (DPO) algorithm (Rafailov et al., 2023) to implicitly align low-step inference trajectories with high-step inference trajectories, eliminating the need for reward models and costly external preference data collection. As shown in Figure 1, DA effectively reduces blurriness while TA improves detail preservation under few-step settings.
@@
2 Related Work
2.1 Video Diffusion Model
Driven by advances in large-scale video data and model architectures, video diffusion models have achieved remarkable success. Early approaches (Blattmann et al., 2023; Guo et al., 2024) using 3D U-Net architectures in latent space were limited to generating short, low-quality videos. With the success of Sora (OpenAI, 2024), modern models have shifted toward Diffusion Transformer (DiT) architectures, with open-source implementations like HunyuanVideo (Kong et al., 2025), CongVideoX (Yang et al., 2025), and Wan (Wan et al., 2025). Although these models can produce longer, higher-quality videos, generation remains computationally expensive and time-consuming. To address this limitation, we leverage diffusion distillation methods to accelerate these models without compromising their quality.
2.2 Diffusion Model Distillation
@@
Trajectory-preserving distillation aims to train a student model capable of generating samples in fewer steps by approximating the ordinary differential equation (ODE) sampling trajectory of the teacher model. Progressive Distillation (Salimans and Ho, 2022) trains the student model to predict the subsequent flow locations determined by the teacher model. Rectified Flow (Liu et al., 2022) trains the student model on noisy-image pairs obtained from the teacher, thereby progressively straightening the ODE trajectory. Consistency Distillation (Luo et al., 2023) trains a student model to map noisy samples along the ODE trajectory to its origin. Although trajectory-preserving distillation facilitates training, the samples generated in few-step tend to exhibit blurriness and inferior quality.
Distribution-matching distillation trains a student model to approximate the teacher’s distribution. Some methods leverage adversarial loss functions to improve the generative quality of student models (Sauer et al., 2024b, a). DMD (Yin et al., 2024b) minimizes approximate KL divergence using gradients from two score functions: a fixed pretrained target score and a dynamically trained generator score.
Recently, trajectory-preserving distillation methods (Li et al., 2024; Wang et al., 2023) and distribution-matching distillation(Zhang et al., 2024)) have been increasingly applied to video generative models. However, many of these distilled video models originate from the pre-Sora era and typically generate only low-quality, short-duration results. Moreover, existing consistency distillation methods for video are trained using discretized timesteps, which rely on numerical ODE solvers and introduce additional discretization errors. In contrast, our approach employs continuous-time consistency distillation to address these issues. Furthermore, prior distribution-matching distillation methods usually approximate the teacher’s output distribution as the target, which implicitly limits the generation quality by the teacher model. Instead, our approach directly performs distribution alignment on real data, similar to (Lin et al., 2025).
@@
3 Preliminary
3.1 Flow Matching
Flow Matching (Lipman et al., 2023; Liu et al., 2022) is widely used in recent advanced video-generation models (Wan et al., 2025). Given a data sample $\mathbf{x}_{0}\sim p_{\textnormal{data}}$ and a noise sample $\mathbf{x}_{1}\sim p_{\textnormal{noise}}$, the noisy data $\mathbf{x}_{t}$ is defined as a linear interpolation between these two distributions:
@@ 3.2.2 Continuous-time CMs
where $\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\mathbf{x}_{t},t)}{\mathrm{d}t}=\nabla_{x_{t}}\bm{f}_{\theta^{-}}(\mathbf{x}_{t},t)\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}+\partial_{t}\bm{f}_{\theta^{-}}(\mathbf{x}_{t},t)$ is the tangent of $\bm{f}_{\theta^{-}}$ at $(\mathbf{x}_{t},t)$ along the trajectory of the PF-ODE $\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$. Continuous-time CMs do not rely on ODE solvers, which avoids discretization errors and offers more accurate supervision signals during training. Recently, sCM (Lu and Song, 2025) presents several training tricks to simplify, stabilize, and scale the training of continuous-time consistency models, achieving superior performance compared to discrete-time consistency models.
4 Method
@@
4.1 Continuous-time Consistency Distillation
@@
$\bm{F}_{\theta}(x,t)$ is a deep neural network. Unlike the TrigFlow formulation employed in sCM (Lu and Song, 2025), we set $c_{\mathrm{skip}}=1$ and $c_{\mathrm{out}}=-t$, which aligns with the Rectified Flow (Liu et al., 2022) formulation commonly adopted in current video diffusion models and eliminates the need for additional formulation conversion. Moreover, this parameterization satisfies the boundary conditions and is compatible with the Euler ODE solver. Then we differentiate both sides of Eq (4) with respect to $t$:
@@
4.1.1 Training Stabilization and Acceleration
Building upon the approach proposed in sCM (Lu and Song, 2025), we employed two training techniques, Tangent Normalization and Tangent Warmup, to mitigate the instability issues in the continuous-time consistency distillation process. Additionally, given that current video diffusion models comprise multiple DiT blocks, we designed an iterative algorithm to compute the tangent function $\frac{\mathrm{d}\bm{f}_{\theta^{-}}(\mathbf{x}_{t},t)}{\mathrm{d}t}$ with minimal computational overhead.
@@ 4.1.1 Training Stabilization and Acceleration
4.2 Alignment for Distilled Model
@@
4.2.1 Distribution Alignment
While continuous-time consistency distillation achieves precise ODE trajectory preservation, the generated results still exhibit blurriness due to the inherent properties of consistency models (Lin et al., 2024). Therefore, we introduce distribution alignment to improve the visual fidelity and realism of generated videos. Unlike previous distribution-matching methods that constrain the upper limit of the student’s generation quality by matching the teacher’s distribution, we directly employ adversarial loss to align with real data distribution.
The distribution alignment employs a discriminator trained to distinguish between generated samples $\hat{\mathbf{x}}_{0}$ and real data $\mathbf{x}_{0}$, where $\hat{\mathbf{x}}_{0}=\mathbf{x}_{t}-t\bm{F}_{\theta}(\mathbf{x}_{t},t)$. For the discriminator, we follow the design in StyleGAN-T (Sauer et al., 2023). We utilize a frozen pretrained feature network DINOv2 (Oquab et al., 2024) coupled with trainable lightweight discriminator heads, comprising both temporal heads for capturing inter-frame temporal dynamics and spatial heads for evaluating intra-frame visual quality.
@@ 4.2.2 Trajectory Alignment
Subsequently, we finetune our model using Direct Preference Optimization (Liu et al., 2025) on this synthetic preference dataset. As shown in Figure 3, since the synthetic data implicitly contains flow trajectory information, this training enables the model to align with high-step inference trajectories, thereby enhancing low-step generation quality. The DPO loss $\mathcal{L}_{\textnormal{DPO}}$ on flow-based video generation model is written below:
@@
Employing Eq. (11) alone as the training objective for trajectory alignment results in undesired training dynamics. As training progresses, the model minimizes the loss by reducing velocity prediction accuracy on dispreferred data more than on preferred data, thereby causing the flow to diverge from the intended trajectory. To address this issue, we introduce a regularization term $\|\mathbf{v}^{w}-\mathbf{v}_{\theta}(\mathbf{x}_{t}^{w},t)\|^{2}$ to maintain high prediction fidelity on the preferred data. Given that our preference dataset is entirely synthetic, this regularization term takes the same form as the reflow loss (Liu et al., 2022), denoted as $\mathcal{L}_{RF}$. Therefore, the overall trajectory alignment loss can be formulated as $\mathcal{L}_{\textnormal{DPO}}+\lambda_{\textnormal{RF}}\cdot\mathcal{L}_{\textnormal{RF}}$, where $\lambda_{\textnormal{RF}}$ is a hyper-parameter. We note that the model performance is robust to the choice of $\lambda_{\textnormal{RF}}$, as detailed in the appendix.
5 Experiment
5.1 Implementation Details.
@@ 5 Experiment
In our distillation experiments, we use Wan2.1-FUN-inp-480p-1.3B (Wan et al., 2025) as the foundational model. For fair comparison, all the experiments are conducted on OpenVid-1M (Nan et al., 2025), a large-scale, high-quality video dataset that provides a standardized benchmark for few-step video generation. Specifically, we randomly select one million videos for training and 1000 for evaluation. In the continuous-time consistency distillation and distribution alignment, we fix the resolution of the training videos to $832\times 480$ with 61 frames and a batch size of 1. We conduct full-parameter training for 3,000 iterations using a learning rate of $1\mathrm{e}-6$ for both the discriminator and student model. The default ema rate is 0.95, warm-up iteration $\mathcal{N}_{\textnormal{warmup}}$ is 1000 and $\lambda_{\textnormal{adv}}$ is set to 0.01. In trajectory alignment, we first utilize distribution-aligned distilled model to generate videos using 4 steps and 8 steps sampling respectively, then construct preference dataset of size 5000. Then we apply LoRA (Hu et al., 2022) to fine-tune the model’s linear layers on this preference dataset, using a learning rate of $1\mathrm{e}-6$ for approximately 2000 iterations. To further improve 2-step generation quality, we conduct an additional round of trajectory alignment, employing the once-aligned model to generate preference data using 2-step and 4-step sampling with the same training procedure. $\lambda_{RF}$ is set to 2 and $\beta$ in $\mathcal{L}_{\mathrm{DPO}}$ is set to 2500. All training uses the AdamW optimizer (Loshchilov and Hutter, 2019), and we employ the Euler solver for sampling.
5.2 Evaluation Metrics
@@
Following existing work (Zhang et al., 2024; Yin et al., 2025), we use FVD (Ge et al., 2024) and standard VBench-I2V benchmark (Huang et al., 2024) to evaluate our model. FVD measures the Fréchet distance between feature distributions of real and generated videos. VBench is a comprehensive benchmark that is widely used to evaluate the video quality. All evaluations are conducted on the OpenVid-1M test set.
5.3 Comparisons on Image-to-Video Generation
5.3.1 Quantitative Comparisons.
@@ 5.3 Comparisons on Image-to-Video Generation
We conduct comprehensive evaluations by comparing our model against state-of-the-art distillation methods, including trajectory-preserving methods LCM (Luo et al., 2023) and PCM (Wang et al., 2024), and distribution-matching methods OSV (Mao et al., 2025) and DMD2 (Yin et al., 2024a). We implement these methods on Wan 2.1 following the official codebases of LCM, PCM, and DMD2. And we try our best to implement OSV, which we denote as $\textnormal{OSV}^{\dagger}$. During evaluation, we utilize the first frame from the Openvid-1M test set as conditional input to generate 1000 videos at 832×480 resolution with 24 FPS, then compute FVD and VBench metrics on these generated videos.
@@
5.3.2 Qualitative Comparisons
Figure 4 presents visual comparison results of the teacher model with 8 sampling steps, LCM (Luo et al., 2023), PCM (Wang et al., 2024), OSV (Mao et al., 2025), DMD2 (Yin et al., 2024a) and our method with 4 inference steps. Figure 4 demonstrates that our method generates videos with superior visual clarity, richer details, and enhanced realism compared to other approaches, while exhibiting fewer artifacts. Additional qualitative results are provided in the appendix.
To further evaluate the performance of our method, we conduct a comprehensive user study comparing it against existing distillation approaches and the teacher model. We select first frames from 50 videos across different categories in the OpenVid-1M test set as conditional inputs. For each image, we generate 5 videos using same random seeds, with distillation methods using 4 sampling steps and the teacher model using 8 steps. Participants evaluate and select the best video based on frame quality, temporal consistency, and image alignment. As shown in Figure 5(a), our approach significantly outperforms baseline distillation methods and the teacher model with 8 sampling steps.
Figure 4: Qualitative comparison results. Our results significantly outperform other methods in both frame quality and temporal consistency under 4 sampling steps. All visual results are generated using images from the OpenVid-1M (Nan et al., 2025) test set.
(a) Results from our user study regarding frame quality, temporal consistency, and image alignment of different models. The teacher method uses 8 sampling steps, while other methods use 4 steps.
(b) Effect of synthetic preference dataset and reflow loss in trajectory alignment.